### v1

# Influence of Non-extensivity on the drag and diffusion coefficients of hadronic matter

Abstract
In this work, we investigate the drag and diffusion coefficients of various hadrons propagating through a hadronic thermal bath by employing the Fokker–Planck equation within the framework of Tsallis non-extensive statistics. The non-extensive parameter $q$ accounts for the deviation from equilibrium and provides a more realistic description of the medium that is not perfectly thermalized. The hadronic bath, consisting of various mesonic and baryonic species, is characterized by different mass cutoffs that control the spectral composition of the medium. Our analysis shows that both the drag $F$ and momentum diffusion coefficients $\Gamma$ increases exponentially with temperature and increases systematically with increasing $q$ and mass cutoff. The spatial diffusion coefficient $D_{x}$ exhibits a decreasing trend with temperature $T$, $q$ and mass cutoff which highlights the significant influence of non-equilibrium effects and hadronic composition on the transport behaviour of hadrons, offering valuable insights into the thermal and dynamical properties of the hadronic phase preceding freeze-out in heavy-ion collisions. Additionally, we have studied the relaxation time of heavy mesons such as $D_{0}$, $J/\psi$ and $\Upsilon$. We found that the heavier mesons relaxed later in comparison to the lighter mesons in the hadronic medium.
pacs:
25.75.-q,25.75.Nq,25.75.Gz, 25.75.Dw,12.38.Mh, 24.85.+p
I Introduction
Relativistic heavy-ion collision experiments at major facilities such as RHIC and the LHC offer an exceptional opportunity to explore the properties of strongly interacting matter at extreme temperatures and densities. In these high-energy environments, ordinary nuclear matter undergoes a remarkable transformation from a confined hadronic phase to a deconfined state composed of quarks and gluons, known as the quark–gluon plasma (QGP) Niida and Miake (2021); Iancu (2014); Steinberg (2011); Wyslouch (2011). This transition marks a fundamental change in the behaviour of matter governed by Quantum Chromodynamics (QCD). Following the collision, the initially formed hot and dense QGP rapidly expands and cools, eventually hadronizing into a system dominated by interacting hadrons. During this hadronic phase, the medium remains strongly coupled, with significant interactions among its constituents before the system reaches thermal and chemical freeze-out He et al. (2012); Cao et al. (2015); Becattini et al. (2014). A comprehensive understanding of this evolution—from the early deconfined stage to the late hadronic phase—requires detailed knowledge of the transport properties such as drag, diffusion and viscosity. These microscopic quantities play a crucial role in determining the rate of thermalization, collective flow and energy loss mechanisms within the medium. By studying these properties, researchers can better characterize the dynamics of the QGP, the nature of the phase transition, and the emergent behaviour of strongly interacting matter created in relativistic heavy-ion collisions.
Heavy mesons are excellent probes for investigating the quark–gluon plasma (QGP) due to their distinct production and interaction characteristics. Heavy quarks, such as charm and bottom are produced in the initial hard scatterings of these collisions, well before the QGP is formed, allowing them to experience the entire evolution of the medium — from the deconfined QGP phase through hadronization to the hadronic stage. Their large masses suppress thermal production, ensuring that any observed modifications in their spectra or flow patterns arise solely from interactions with the medium rather than from new production processes.Cao et al. (2015); He et al. (2012); Svetitsky (1988) The propagation of heavy quarks through the QGP involves numerous soft scatterings with medium constituents, leading to momentum degradation (drag) and momentum broadening (diffusion), which can be effectively described by the Fokker–Planck transport approach. These transport coefficients directly encode the strength of the coupling between heavy quarks and the QGP. Furthermore, the resonant interactions near the critical temperature $T_{c}$ enhance the coupling between heavy quarks and the medium, resulting in stronger drag and diffusion. During hadronization, heavy quarks recombine with light quarks through coalescence processes that retain imprints of the medium’s dynamics. Consequently, the study of heavy-meson transport and diffusion coefficients provides a sensitive means of characterizing the coupling strength, microscopic dynamics, and transport properties of both the quark–gluon plasma and the hadronic matter phases.
The drag and diffusion coefficients constitute fundamental transport parameters that characterize the interaction of a propagating particle with a thermalized medium. The drag coefficient quantifies the average rate of momentum loss due to dissipative forces arising from successive scatterings with the medium constituents. While the momentum diffusion coefficient describes the stochastic broadening of the particle’s momentum distribution, reflecting the random fluctuations induced by microscopic collisions. On the other hand, spatial diffusion coefficient governs the rate at which the particle spreads through space and is inversely related to the strength of its coupling with the medium. Together, these coefficients provide a comprehensive description of the dissipative and fluctuating dynamics of hadrons, determining their relaxation and equilibration time scales and thereby offering crucial insights into energy loss mechanisms and transport properties within the strongly interacting matter produced in relativistic heavy-ion collisions.Du (2024); Madni et al. (2023)
To analyze such transport processes in a thermalized medium, it is essential to employ the Fokker–Planck equation, which offers a robust and systematic framework for describing particle dynamics in momentum space. This equation serves as a crucial link between the microscopic scattering interactions of particles and their macroscopic transport behaviour Lang et al. (2016); van Hees et al. (2008); Svetitsky (1988); Golam Mustafa et al. (1998); van Hees and Rapp (2005). In the context of hadronic medium, a propagating hadron undergoes numerous soft collisions with the surrounding constituents. These repeated stochastic interactions lead to continuous changes in its momentum, effectively characterized by drag (systematic loss of momentum) and momentum diffusion (random momentum broadening). The Fokker–Planck approach encapsulates these effects through transport coefficients that quantify the rate of momentum relaxation and the extent of thermal fluctuations. Through this formalism, one can trace how a particle’s momentum distribution evolves toward equilibrium, offering insights into the medium’s coupling strength and its dynamical properties. Moreover, it enables the study of equilibration timescales, energy loss mechanisms, and the degree of thermalization experienced by the propagating particle.
In the small-momentum-transfer regime, where individual collisions impart only minor perturbations to the particle’s total momentum, the Boltzmann transport equation can be systematically reduced to the Fokker–Planck formAndronic and others (2016). This reduced form provides a unified kinetic description incorporating both drag and diffusion components, making it particularly well suited for investigating the dynamical evolution, energy dissipation and transport behaviour of hadrons in hot and dense hadronic or partonic environments.
Recent studies have utilized a range of QCD-inspired theoretical approaches to investigate the temperature dependence of drag and diffusion coefficients for hadrons and heavy quarks. In reference Laine (2011), Heavy Meson Chiral Perturbation Theory was employed to analyze diffusion coefficient, while reference He et al. (2011) examined transport coefficients using empirical elastic scattering amplitudes of $D$ mesons with thermal hadrons. Additionally, effective field theory–based calculations of drag and diffusion were carried out in references Ghosh et al. (2011); van Hees and Rapp (2005) and reference He et al. (2011) utilized vacuum scattering amplitudes derived from effective Lagrangians. More recently, transport properties have also been explored within the Color String Percolation Model (CSPM) Goswami et al. (2023b) and the Van der Waals Hadron Resonance Gas (VDWHRG) framework Goswami et al. (2023a). Traditionally, these studies are carried out assuming a medium in perfect thermal equilibrium, described by Boltzmann–Gibbs statistics. However, experimental observations indicate that the particle momentum spectra in heavy-ion collisions often exhibit power-law tails, suggesting significant non-equilibrium effects. To capture such deviations, the Tsallis non-extensive statistics has emerged as a powerful generalization of conventional thermodynamicsAkhil and Tiwari (2024); Singh et al. (2024). The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium, with $q=1$ recovering the Boltzmann–Gibbs limit and $q>1$ representing systems with long-range correlations and memory effectsTsallis (1988); Kyan and Monnai (2022); Zheng et al. (2015); Singh et al. (2024). In this framework, the non-extensive parameter $q$ modifies the particle distribution functions, directly affecting the drag and diffusion coefficients. Hence, their study under Tsallis statistics reveals key insights into the non-equilibrium transport behaviour of hadrons in a thermal hadronic medium.
The manuscript is organized as follows: In Section II we outline the process of determining the drag and diffusion coefficients using the Fokker-Planck equation within the Tsallis distribution function. In Section III, we have done a comprehensive discussion of the results obtained. Lastly, in Section IV, we provide a succinct summary of the study.
II Formulation
II.1 The Drag and Diffusion of Hadron gas
To investigate the transport dynamics of hadrons in QGP produced in relativistic heavy-ion collisions, we employ the Fokker–Planck equation as a theoretical framework which provides a microscopic description of the stochastic evolution of the hadrons momentum distribution due to its interaction with the thermalized medium constituents which is light quarks. The dissipative and fluctuating aspects of the medium-induced transport which is explained by drag and diffusion coefficients, play a central role in quantifying the degree of thermalization and energy loss mechanisms of hadrons within the QGP.
The Fokker-Planck equation in three dimension is given by Abreu et al. (2011)
$\frac{\partial f_{c}}{\partial t}={F}\vec{\nabla_{p}}.(\vec{P}f_{c})+\Gamma\nabla^{2}_{p}f_{c}$ (1)
Here $F$, $\Gamma_{0}$, $\Gamma_{1}$ are scalar coefficients and do not depends on $\vec{P}$ which means $F(P^{2})=F$ and $\Gamma_{0}(P^{2})$= $\Gamma_{1}(P^{2})$. The first term on the right-hand side corresponds to the drag force, representing the systematic loss of momentum due to interactions with the surrounding. The second term involves the momentum diffusion, characterizing stochastic momentum broadening caused by random scattering with medium constituents.
The standard diffusion equation is given by the equation
$\frac{\partial C}{\partial t}=-{\mu}\nabla.(C\vec{F})+D\nabla^{2}C$ (2)
where $\mu$= mobility, C = concentration of a solute, D= diffusion coefficient, $\vec{F}$= external force.
On comparing the above vector equations we get F as friction coefficient or drag and $\Gamma$ as diffusion coefficient which characterize the dissipative and fluctuating interactions with the medium.
The F and $\Gamma$} are the moment of transition rates in Fokker-Planck equation 1 and is given by Ozvenchuk et al. (2014)-
$F_{i}(\vec{P})=\int K_{i}w(\vec{P},\vec{K})\,d\vec{K}$ (3)
The momentum-space diffusion matrix which accounts for broadening of momentum distribution is:
$\Gamma_{ij}(\vec{P})=\frac{1}{2}\int K_{i}K_{j}w(\vec{P},\vec{K})\,d\vec{K},$ (4)
where $w(\vec{P},\vec{K})$} is transition rate or collision rate, $\vec{P}$ is initial momenta, $\vec{K}$ is transferred momenta, $\vec{P}-\vec{K}$ is final momenta.Tolos et al. (2016)
In the ideal scenario of a homogeneous and isotropic pion gas, the coefficients $F_{i}$ and $\Gamma_{ij}$ which depends solely on the particle momentum $\vec{P}(P_{i})$, can be conveniently expressed in terms of three independent scalar functions $F$, $\Gamma_{0}$ and $\Gamma_{1}$ as $F_{i}(\vec{P})=F(P^{2})P_{i}$ and $\Gamma_{ij}(\vec{P})=\Gamma_{0}(P^{2})\Delta_{ij}+\Gamma_{1}(P^{2})\frac{P_{i}P_{j}}{P^{2}}$ where $\Delta_{ij}=\delta_{ij}-\frac{P_{i}P_{j}}{P^{2}}$ and $\Delta_{ij}\Delta^{ij}=2$ Abreu et al. (2011).
Now Fokker Planck in one dimension is given by,
$\frac{\partial f_{c}}{\partial t}={F}\frac{\partial(pf_{c})}{\partial p}+\Gamma\frac{\partial^{2}f_{c}}{\partial p^{2}}$ (5)
The above equation also known as Rayleigh’s equation which describes momentum distributon equation for a brownian particle.
Now using initial conditions ${f_{c}}(p,t=0)=\delta(p-p_{0})$ The analytic solution of equation 5 is given by Abreu et al. (2011),
$f_{c}(p,t)=\left[\frac{F}{2\pi\Gamma}(1-e^{-2Ft})\right]^{\frac{-1}{2}}\exp{\frac{-F(p-p_{0}e^{-Ft})^{2}}{2\Gamma(1-e^{-2Ft})}}$ (6)
The term $p>=p_{0}e^{-Ft}$ represents the exponential decay of the mean momentum of the particle due to the action of the drag force, which continuously dissipates its initial momentum through successive interactions with the surrounding medium. Simultaneously, the factor $p^{2}>-<p>^{2}=\frac{\Gamma}{F}(1-e^{-2Ft})$ in the denominator governs the broadening of the momentum distribution.
Here F and $\Gamma$} are not independent but related by Einstein’s fluctuation-dissipation relation
$\Gamma=FmT$ (7)
and the momentum diffusion coefficient $\Gamma$ is related to spatial diffusion coefficient $D_{x}$ as
$D_{x}=\frac{\Gamma}{{m^{2}}{F^{2}}}=\frac{T^{2}}{\Gamma}=\frac{T}{{m}{F}}$ (8)
Now if $p(t)$ is survival probability = $e^{-wt}$ where $w$ is transition rate or collision rate that depends on velocity $v$. Then $p(t)dt$ = survival probability for $dt$ time interval = $we^{-wt}dt$. The relaxation time or mean time is $\tau$ = $\bar{t}$ = $\int_{0}^{\infty}t.p(t)dt$ = $\int_{0}^{\infty}wt.e^{-wt}dt$ = $\frac{1}{w}$ OR $\tau^{-1}=w$. For the process $a(p_{a})+b(p_{b})\rightarrow a(p_{c})+b(p_{d})$, if $n_{a}$ is number of particles per unit volume with relative velocity $v_{ab}$ contained in volume $v_{ab}dtdA$ , then relative incident flux = $n_{a}v_{ab}$. Number of scattered particles will give collision probability as Reif (1965)-
$w=\tau^{-1}=\frac{(n_{a}v_{ab})\sigma_{ab}(n_{b}d^{3}r)}{n_{a}d^{3}r}$ (9)
$\tau^{-1}=n_{b}v_{ab}\sigma_{ab}$ (10)
Average relaxation time
$\tilde{\tau_{a}}^{-1}=\frac{\int{\frac{d^{3}p_{a}}{(2\pi)^{3}}\tau^{-1}f_{a}^{0}}}{\int{\frac{d^{3}p_{a}}{(2\pi)^{3}}f_{a}^{0}}}$ (11)
$\tilde{\tau_{a}}^{-1}=\sum_{b}\frac{\int{\frac{d^{3}p_{a}}{(2\pi)^{3}}\frac{d^{3}p_{b}}{(2\pi)^{3}}\sigma_{ab}v_{ab}f_{a}^{0}f_{b}^{0}}}{\int{\frac{d^{3}p_{a}}{(2\pi)^{3}}f_{a}^{0}}}$ (12)
$\tilde{\tau_{a}}^{-1}=\sum_{b}n_{b}<\sigma_{ab}v_{ab}>$ (13)
Here $n_{b}=\int\frac{d^{3}p_{b}}{(2\pi)^{3}}f_{b}^{0}$ is the number density and $f_{b}^{0}$ is the distribution function of $b^{th}$ hadronic species.
General definition of thermal average
$\sigma_{ab}v_{ab}>=\frac{\int d^{3}p_{a}d^{3}p_{b}e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}\sigma v_{ab}}{d^{3}p_{a}d^{3}p_{b}e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}}$ (14)
Assuming constant cross section, $\sigma$ of the same species at zero baryon density. Here in this work we use the Tsallis distribution (in the local rest frame) given by $f_{a}^{0}=[1-(q-1)\frac{E_{a}}{T}]^{q/(q-1)}=e_{q}^{-E_{a}/T}$.
Now the momentum space volume elements can be written as Tiwari et al. (2018)
$d^{3}p_{a}d^{3}p_{b}=8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta$
and
$v_{ab}=\frac{\sqrt{(E_{a}E_{b}-p_{a}p_{b}\cos\theta)^{2}-(m_{a}m_{b})^{2}}}{E_{a}E_{b}-p_{a}p_{b}\cos\theta}$ (15)
$<\sigma_{ab}v_{ab}>=\frac{\sigma\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}\times\frac{\sqrt{(E_{a}E_{b}-p_{a}p_{b}cos\theta)^{2}-(m_{a}m_{b})^{2}}}{E_{a}E_{b}-p_{a}p_{b}cos\theta}}}{\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}}}$ (16)
The relaxation time is related to drag coefficient as $\tau^{-1}=F$
Thus drag
$F=\tilde{\tau_{a}}^{-1}=\sum_{b}n_{b}<\sigma_{ab}v_{ab}>$ (17)
$F=\sum_{b}n_{b}\frac{\sigma\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}\times\frac{\sqrt{(E_{a}E_{b}-p_{a}p_{b}cos\theta)^{2}-(m_{a}m_{b})^{2}}}{E_{a}E_{b}-p_{a}p_{b}cos\theta}}}{\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}}}$ (18)
The momentum diffusion coefficient $\Gamma$ is given as
$\Gamma=FmT=mT\sum_{b}n_{b}<\sigma_{ab}v_{ab}>$ (19)
$\Gamma=mT\sum_{b}n_{b}\frac{\sigma\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}\times\frac{\sqrt{(E_{a}E_{b}-p_{a}p_{b}cos\theta)^{2}-(m_{a}m_{b})^{2}}}{E_{a}E_{b}-p_{a}p_{b}cos\theta}}}{\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}}}$ (20)
The spatial diffusion coefficient $D_{x}$ is given as
$D_{x}=\frac{T}{mF}=\frac{T}{m\sum_{b}n_{b}<\sigma_{ab}v_{ab}>}$ (21)
$D_{x}=\frac{T}{\displaystyle{m\sum_{b}n_{b}\frac{\sigma\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}\times\frac{\sqrt{(E_{a}E_{b}-p_{a}p_{b}cos\theta)^{2}-(m_{a}m_{b})^{2}}}{E_{a}E_{b}-p_{a}p_{b}cos\theta}}}{\int{8\pi^{2}p_{a}p_{b}dE_{a}dE_{b}dcos\theta e_{q}^{-E_{a}/T}e_{q}^{-E_{b}/T}}}}}$ (22)
(a) Variation of Drag coefficient with temperature for various q-values
(b) Variation of Drag coefficient with temperature for various mass cutoffs
Figure 1: Variation of Drag coefficient of hadron gas as a function of temperature for various q values and mass cutoffs.
III Results and Discussions
In this section, we present a detailed analysis of the drag and diffusion coefficients of hadrons propagating through a hadronic thermal medium. The transport coefficients have been evaluated using the Fokker–Planck formalism within the framework of non-extensive Tsallis statistics, which provides a generalized description of systems that deviate from ideal equilibrium conditions. The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium and serves as a crucial probe of the medium’s thermodynamic state, offering a more realistic representation of the hadronic environment formed in relativistic heavy-ion collisions where complete thermalization is rarely achieved. The interactions of hadrons with the surrounding light and heavy quarks constitute the effective thermal bath, whose microscopic properties are influenced by the inclusion of higher resonant statesvan Hees and Rapp (2005). To account for these contributions, different hadronic mass cutoffs have been implemented, allowing a systematic study of how heavier resonances modify the thermal density, scattering rates and overall transport behaviour of the medium. The mass cutoff thus plays a significant role in determining the effective strength of interactions between the probe particle and the medium constituentsCastorina et al. (2010). The variation of the drag and diffusion coefficients with the Tsallis parameter $q$ reveals the sensitivity of the transport properties to non-equilibrium effects in the medium. Thus, the Tsallis non-extensive parameter $q$ and the chosen mass cutoff provides important insights into the interplay between thermal and non-equilibrium effects in the strongly interacting hadronic phase.
In Figure 1, we have shown the temperature dependence of the drag coefficient for various values of the Tsallis non-extensivity parameter $q$ by using Eq.(18). We notice the drag coefficient increases exponentially with temperature and a systematic increase in drag while increase of non-extensive parameter $q$ and mass cutoffs which contitute the thermal bath.
The drag coefficient exhibits an exponential rise with increasing temperature, reflecting the growing strength of interactions between the propagating hadron and the constituents of the hadronic medium. As the temperature increases, the thermal population and average momentum of hadrons become significantly higher, leading to an enhanced rate of collisions and stronger momentum dissipation. This behaviour signifies that the medium becomes increasingly dense and interactive as it approaches the critical temperature Bazavov and others (2014), indicating the onset of strong coupling in the hadronic phase.
(a) Variation of Drag coefficient of $D_{0}$ with temperature for various mass cutoffs
(b) Variation of Drag coefficient of $J/\psi$ with temperature for various mass cutoffs
(c) Variation of Drag coefficient of $\Upsilon$ with temperature for various mass cutoffs.
Figure 2: Variation of Drag coefficient of heavy mesons as a function of temperature for various mass cutoffs.
(a) Relaxation times of various heavy mesons versus temperature.
(b) Ratios of relaxation time of $J/\psi$ over $D_{0}$ and $\Upsilon$ over $D_{0}$.
Figure 3: Relaxation times of $D_{0}$, $J/\psi$ and $\Upsilon$ and their ratios.
A clear enhancement of the drag coefficient is observed with the increase of the Tsallis non-extensive parameter $q$ as seen in fig 1(a). This parameter quantifies the deviation of the system from thermal equilibrium, and values of $q>1$ correspond to a medium with a longer high-energy tail in its momentum distribution. Such non-equilibrium effects enhance the probability of energetic scatterings between the probe and medium constituents, resulting in higher momentum transfer and consequently larger drag. The systematic increase of the drag coefficient with $q$ therefore reflects the growing influence of non-equilibrium dynamics on the transport behaviour of the hadronic medium.
The drag coefficient also increases with the hadronic mass cutoff as seen in fig 1(b), which determines the extent to which heavier resonant states are included in the thermal spectrum. A higher cutoff incorporates additional massive resonances, thereby increasing the effective density and the number of available scattering channels in the medium. The inclusion of these heavy states enhances the overall interaction strength between the heavy resonant states and the surrounding hadrons, leading to a noticeable rise in the drag coefficient. This trend underscores the role of hadronic resonances in strengthening transport interactions and modifying the microscopic dynamics of the thermal bath near critical temperature Bazavov and others (2014).
(a) Variation of Momentum Diffusion coefficient of $D_{0}$ meson with temperature for various q-values.
(b) Variation of Momentum Diffusion coefficient of $D_{0}$ meson with temperature for various mass cutoffs.
Figure 4: Variation of Momentum Diffusion coefficient of $D_{0}$ meson with temperature for various q values and mass cutoffs.
Figure 2 depicts the drag coefficient for various heavy mesons $D_{0}$, $J/\psi$ and $\Upsilon$ propagating through a thermal bath of hadronic medium controlled by mass cutoffs. The drag coefficient, which represents the thermal average of the squared momentum transfer between the heavy mesons and the thermal bath particles, shows a clear temperature dependence. As the temperature of the thermal bath increases, the kinetic energy of the constituent hadrons also rises, enhancing their capability to transfer larger momentum during interactions with the heavy mesons. Consequently, this leads to an increase in the drag coefficient with temperature as observed in Fig  2. We also notice drag of $D_{0}$ meson is greater than drag of $J/\psi$ and $\Upsilon$ and drag of $\Upsilon$ is lowest among them. It can be explained by the expression of drag coefficient in nonrelativistic regime $F\propto\frac{1}{M}$ Abreu et al. (2013), the inverse dependence of drag on the heavy mesonic mass which makes the drag coefficient smaller for heavier mesons which means larger mesonic mass suppresses momentum transfer in medium interactions, leading to a weaker drag coefficient Torres-Rincon et al. (2016); Tolos and Torres-Rincon (2013).
The relaxation time $\tau$ = $F^{-1}$ of heavy mesons is presented in figure 3 together with their corresponding ratios. A clear increase in the relaxation time is observed with increasing mesons mass, indicating that heavier mesons require a longer time to relax in the hadronic medium. Consequently, the heavier mesons are unlikely to fully relax before exiting the hadronic medium. Furthermore, the ratios of relaxation time remains nearly constant over the considered temperature range. Quantitatively, the relaxation-time ratio lies in the range 2.3-2.9 for $J/\psi$ relative to $D_{0}$ and a ratio of 1.6-1.9 is obtained for $\Upsilon$ relative to $D_{0}$.
Figure 4 illustrates the variation of the momentum diffusion coefficient of the $D_{0}$ meson with temperature for different values of the Tsallis parameter $q$ and hadronic mass cutoffs using Eq.(20). We notice an exponential increase in momentum diffusion along with temperature align with Ghosh et al. (2011); Torres-Rincon et al. (2022) and a systematic increase with parameter $q$ and mass cutoffs. The coefficient exhibits a pronounced exponential rise with increasing temperature, reflecting the enhanced random momentum transfer between mesons and the surrounding hadrons at higher thermal energies. As the medium becomes hotter and denser, the frequency and strength of particle collisions increases, amplifying stochastic interactions and momentum fluctuations. This behaviour indicates a transition towards a more strongly coupled regime near the critical temperature, where thermal agitation and scattering effect dominates the transport dynamics. The momentum diffusion coefficient also increases systematically with the Tsallis non-extensive parameter $q$ can be seen in fig 4(a). A higher value of $q$ represents stronger deviation from equilibrium and a broader momentum distribution within the medium. This extended high-energy tail enhances the probability of large-momentum transfer processes, thereby intensifying the random kicks imparted to the propagating mesons. A similar increasing trend is observed with the hadronic mass cutoff in fig 4(b) which defines the upper limit of resonant states contributing to the thermal bath. As the mass cutoff increases, more massive hadronic resonances are included, enriching the medium’s particle content and interaction strength. The presence of the heavier states provides additional scattering channels and enhances the overall collision rate, leading to greater momentum diffusion. This dependence highlights the significant role of the hadronic spectrum in governing the transport behaviour of hadrons and further supports the picture of a dense and strongly interacting medium.
(a) Variation of Spatial Diffusion coefficient of $D_{0}$ with temperature for various q-values.
(b) Variation of Spatial Diffusion coefficient of $D_{0}$ with temperature for various mass cutoffs.
Figure 5: Variation of Spatial Diffusion coefficient with temperature of $D_{0}$ meson.
Figure 5 shows the temperature dependence of the spatial diffusion coefficient of the $D_{0}$ meson for different values of the Tsallis parameter $q$ and hadronic mass cutoffs using Eq.(22). A consistent decreasing trend is observed with increasing temperature, $q$, and mass cutoff similarly observed in references He et al. (2011); Torres-Rincon et al. (2022); Ozvenchuk et al. (2014). This behaviour is consistent with the corresponding rise in the drag and momentum diffusion coefficients, since the spatial diffusion coefficient $D_{x}$ is inversely related to these quantities (typically $D_{x}$$\frac{T^{2}}{\Gamma}$$\frac{T}{mF}$). As the temperature rises, the hadronic medium becomes denser and more interactive, resulting in stronger scattering and shorter mean free paths for the propagating mesons. Consequently, the suppression of $D_{x}$ with temperature indicates stronger coupling between mesons and the medium. The spatial diffusion coefficient also decreases with increasing Tsallis parameter $q$. A higher $q$ value reflects a stronger deviation from equilibrium and the emergence of high-energy tails in the particle momentum distribution. These non-equilibrium effects intensify the collision frequency and the effective momentum transfer between the probe and the thermal bath, leading to reduced spatial diffusion. In other terms, an increase in $q$ enhances the medium’s ability to trap or thermalize the propagating particle, thereby lowering its spatial diffusion coefficient and signaling the development of a more strongly correlated, non-equilibrated hadronic environment. A similar decreasing trend of $D_{x}$ is observed with increasing hadronic mass cutoff as in fig 5(b). The inclusion of heavier resonances through higher cutoffs enriches the hadronic spectrum, increasing the effective density and interaction strength of the medium. This enhancement in scattering processes reduces the mean free path of the hadrons, leading to a lower spatial diffusion coefficient. The suppression of $D_{x}$ with the mass cutoff thus emphasizes the critical role of heavy hadronic states in determining transport properties and supports the picture of a dense, strongly interacting medium.
IV Summary and Outlook
In summary, the drag and diffusion coefficients of hadrons have been investigated within the Fokker–Planck framework using Tsallis non-extensive statistics, accounting for the non-equilibrium nature of the hadronic medium produced in relativistic heavy-ion collisions.
  1. 1.
    The drag coefficient exhibits an exponential rise with temperature, suggesting that the interaction between the probe particle and the medium becomes increasingly stronger at higher temperatures. This behaviour reflects the enhanced coupling strength of the system. Such increase signifies more frequent and intense scattering processes within the quark-gluon plasma.
  2. 2.
    The drag coefficient also shows an increasing trend with the non-extensive parameter $q$, indicating the significant role of non-equilibrium dynamics in the system. As $q$ deviates from unity, the medium departs from thermal equilibrium, leading to enhanced momentum transfer between the particles. This rise with $q$ highlights how non-extensivity contribute to greater momentum dissipation and stronger medium interactions.
  3. 3.
    An increase in the hadronic mass cutoff leads to a higher drag coefficient, as heavier resonances are incorporated into the medium. The presence of these massive states enhances the overall interaction rate and scattering processes. Consequently, the medium becomes denser and more strongly coupled, resulting in greater resistance to particle motion and increased drag.
  4. 4.
    The momentum diffusion coefficient also exhibits a similar rising trend with temperature, the non-extensive parameter $q$, and the hadronic mass cutoff. This trend indicates intensified random momentum exchanges between particles and the medium. The combined effects of higher temperature, stronger non-equilibrium dynamics, and heavier resonances contribute to enhanced stochastic momentum transfer within the system.
  5. 5.
    In contrast to the drag and momentum diffusion coefficients, the spatial diffusion coefficient decreases with increasing temperature, non-extensive parameter $q$, and hadronic mass cutoff. This decline indicates that the particle’s mobility within the medium becomes progressively limited under such conditions. The stronger interactions at higher temperature and with heavier resonances lead to reduced spatial transport. Consequently, the meson experiences enhanced confinement and a higher degree of localization within the quark-gluon medium.
  6. 6.
    The relaxation time increases with heavy meson mass, implying that heavier mesons relax more slowly than lighter ones in the medium. The ratios of relaxation times of heavy mesons remain nearly constant over the considered temperature range, indicating a similar temperature response for heavy mesons considered under this study.
  7. 7.
    These overall trend point towards the formation of a dense and strongly interacting hadronic medium. The enhanced coupling and reduced particle mobility reflects the onset of strong correlations and collective behaviour within the medium.
ACKNOWLEDGEMENTS
AKS is grateful to University Grants Commission (UGC), New Delhi for providing research fellowship. SKT acknowledges the financial support of the seed money grant provided by the University of Allahabad, Prayagraj.

### v2

@@
 Additionally, we have studied the relaxation time of heavy mesons such as $D_{0}$, $J/\psi$ and $\Upsilon$.
-We found that the heavier mesons relaxed later in comparison to the lighter mesons in the hadronic medium.
+We found that the heavier mesons relaxed later in comparison to the lighter mesons in the hadronic medium at a particular temperature.
 pacs:
@@
 (2014).
-A comprehensive understanding of this evolution—from the early deconfined stage to the late hadronic phase—requires detailed knowledge of the transport properties such as drag, diffusion and viscosity.
+A comprehensive understanding of this evolution from the early deconfined stage to the late hadronic phase requires detailed knowledge of the transport properties such as drag, diffusion and viscosity.
 These microscopic quantities play a crucial role in determining the rate of thermalization, collective flow and energy loss mechanisms within the medium.
-By studying these properties, researchers can better characterize the dynamics of the QGP, the nature of the phase transition, and the emergent behaviour of strongly interacting matter created in relativistic heavy-ion collisions.
-Heavy mesons are excellent probes for investigating the quark–gluon plasma (QGP) due to their distinct production and interaction characteristics.
-Heavy quarks, such as charm and bottom are produced in the initial hard scatterings of these collisions, well before the QGP is formed, allowing them to experience the entire evolution of the medium — from the deconfined QGP phase through hadronization to the hadronic stage.
+By studying these properties, researchers can better characterize the dynamics of the medium, the nature of the phase transition, and the emergent behaviour of strongly interacting matter created in relativistic heavy-ion collisions.
+Heavy mesons are excellent probes for investigating the strongly interacting matter due to their distinct production and interaction characteristics.
+Heavy quarks, such as charm and bottom are produced in the initial hard scatterings of these collisions, well before the QGP is formed, allowing them to experience the entire evolution of the medium from the deconfined QGP phase through hadronization to the hadronic stage.
 Their large masses suppress thermal production, ensuring that any observed modifications in their spectra or flow patterns arise solely from interactions with the medium rather than from new production processes.Cao et al.
 (2015); He et al.
-(2012); Svetitsky (1988) The propagation of heavy quarks through the QGP involves numerous soft scatterings with medium constituents, leading to momentum degradation (drag) and momentum broadening (diffusion), which can be effectively described by the Fokker–Planck transport approach.
-These transport coefficients directly encode the strength of the coupling between heavy quarks and the QGP.
-Furthermore, the resonant interactions near the critical temperature $T_{c}$ enhance the coupling between heavy quarks and the medium, resulting in stronger drag and diffusion.
-During hadronization, heavy quarks recombine with light quarks through coalescence processes that retain imprints of the medium’s dynamics.
-Consequently, the study of heavy-meson transport and diffusion coefficients provides a sensitive means of characterizing the coupling strength, microscopic dynamics, and transport properties of both the quark–gluon plasma and the hadronic matter phases.
+(2012); Svetitsky (1988) The propagation of heavy mesons through the hadronic medium involves numerous soft scatterings with medium constituents, leading to momentum degradation (drag) and momentum broadening (diffusion), which can be effectively described by the Fokker–Planck transport approach.
+These transport coefficients directly encode the strength of the coupling between heavy mesons and the hadronic medium.
+Consequently, the study of heavy mesons transport and diffusion coefficients provides a sensitive means of characterizing the coupling strength, microscopic dynamics, and transport properties of the hadronic matter.
 The drag and diffusion coefficients constitute fundamental transport parameters that characterize the interaction of a propagating particle with a thermalized medium.
@@
 This reduced form provides a unified kinetic description incorporating both drag and diffusion components, making it particularly well suited for investigating the dynamical evolution, energy dissipation and transport behaviour of hadrons in hot and dense hadronic or partonic environments.
-Recent studies have utilized a range of QCD-inspired theoretical approaches to investigate the temperature dependence of drag and diffusion coefficients for hadrons and heavy quarks.
+Recent studies have utilized a range of QCD-inspired theoretical approaches to investigate the temperature dependence of drag and diffusion coefficients for hadrons and heavy mesons.
 In reference Laine (2011), Heavy Meson Chiral Perturbation Theory was employed to analyze diffusion coefficient, while reference He et al.
@@
 Additionally, effective field theory–based calculations of drag and diffusion were carried out in references Ghosh et al.
-(2011); van Hees and Rapp (2005) and reference He et al.
-(2011) utilized vacuum scattering amplitudes derived from effective Lagrangians.
+(2011); van Hees and Rapp (2005).
+In reference He et al.
+(2011) vacuum scattering amplitudes derived from effective Lagrangians are utilized to study the drag and diffusion coefficients.
 More recently, transport properties have also been explored within the Color String Percolation Model (CSPM) Goswami et al.
@@
 In this framework, the non-extensive parameter $q$ modifies the particle distribution functions, directly affecting the drag and diffusion coefficients.
-Hence, their study under Tsallis statistics reveals key insights into the non-equilibrium transport behaviour of hadrons in a thermal hadronic medium.
+Hence, the study under Tsallis statistics reveals key insights into the non-equilibrium transport behaviour of hadrons in a thermal hadronic medium.
 The manuscript is organized as follows: In Section II we outline the process of determining the drag and diffusion coefficients using the Fokker-Planck equation within the Tsallis distribution function.
@@
 II.1 The Drag and Diffusion of Hadron gas
-To investigate the transport dynamics of hadrons in QGP produced in relativistic heavy-ion collisions, we employ the Fokker–Planck equation as a theoretical framework which provides a microscopic description of the stochastic evolution of the hadrons momentum distribution due to its interaction with the thermalized medium constituents which is light quarks.
-The dissipative and fluctuating aspects of the medium-induced transport which is explained by drag and diffusion coefficients, play a central role in quantifying the degree of thermalization and energy loss mechanisms of hadrons within the QGP.
-The Fokker-Planck equation in three dimension is given by Abreu et al. (2011)
+To investigate the transport dynamics of hadrons produced in relativistic heavy-ion collisions, we employ the Fokker–Planck equation as a theoretical framework which provides a microscopic description of the stochastic evolution of the hadrons momentum distribution due to its interaction with the thermalized medium constituents.
+The dissipative and fluctuating aspects of the medium induced transport is explained by drag and diffusion coefficients which play a central role in quantifying the degree of thermalization and energy loss mechanisms of hadrons within the medium.
+The Fokker-Planck equation in three dimension is given by Abreu et al.
+(2011),
 $\frac{\partial f_{c}}{\partial t}={F}\vec{\nabla_{p}}.(\vec{P}f_{c})+\Gamma\nabla^{2}_{p}f_{c}$ (1)
@@
 The second term involves the momentum diffusion, characterizing stochastic momentum broadening caused by random scattering with medium constituents.
-The standard diffusion equation is given by the equation
+The standard diffusion equation is given as,
 $\frac{\partial C}{\partial t}=-{\mu}\nabla.(C\vec{F})+D\nabla^{2}C$ (2)
@@
 The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium and serves as a crucial probe of the medium’s thermodynamic state, offering a more realistic representation of the hadronic environment formed in relativistic heavy-ion collisions where complete thermalization is rarely achieved.
-The interactions of hadrons with the surrounding light and heavy quarks constitute the effective thermal bath, whose microscopic properties are influenced by the inclusion of higher resonant statesvan Hees and Rapp (2005).
+The interactions of hadrons with the surrounding medium which constitute the effective thermal bath, whose microscopic properties are influenced by the inclusion of higher resonant statesvan Hees and Rapp (2005).
 To account for these contributions, different hadronic mass cutoffs have been implemented, allowing a systematic study of how heavier resonances modify the thermal density, scattering rates and overall transport behaviour of the medium.
@@
 As the temperature increases, the thermal population and average momentum of hadrons become significantly higher, leading to an enhanced rate of collisions and stronger momentum dissipation.
-This behaviour signifies that the medium becomes increasingly dense and interactive as it approaches the critical temperature Bazavov and others (2014), indicating the onset of strong coupling in the hadronic phase.
 (a) Variation of Drag coefficient of $D_{0}$ with temperature for various mass cutoffs
@@
 The inclusion of these heavy states enhances the overall interaction strength between the heavy resonant states and the surrounding hadrons, leading to a noticeable rise in the drag coefficient.
-This trend underscores the role of hadronic resonances in strengthening transport interactions and modifying the microscopic dynamics of the thermal bath near critical temperature Bazavov and others (2014).
+This trend underscores the role of hadronic resonances in strengthening transport interactions and modifying the microscopic dynamics of the thermal bath.
 (a) Variation of Momentum Diffusion coefficient of $D_{0}$ meson with temperature for various q-values.
@@
 Furthermore, the ratios of relaxation time remains nearly constant over the considered temperature range.
-Quantitatively, the relaxation-time ratio lies in the range 2.3-2.9 for $J/\psi$ relative to $D_{0}$ and a ratio of 1.6-1.9 is obtained for $\Upsilon$ relative to $D_{0}$.
+Quantitatively, the relaxation-time ratio lies in the range 2.3-2.9 for $\Upsilon$ relative to $D_{0}$ and a ratio of 1.6-1.9 is obtained for $J/\psi$ relative to $D_{0}$.
 Figure 4 illustrates the variation of the momentum diffusion coefficient of the $D_{0}$ meson with temperature for different values of the Tsallis parameter $q$ and hadronic mass cutoffs using Eq.(20).
@@
 As the medium becomes hotter and denser, the frequency and strength of particle collisions increases, amplifying stochastic interactions and momentum fluctuations.
-This behaviour indicates a transition towards a more strongly coupled regime near the critical temperature, where thermal agitation and scattering effect dominates the transport dynamics.
 The momentum diffusion coefficient also increases systematically with the Tsallis non-extensive parameter $q$ can be seen in fig 4(a).
@@
 This behaviour reflects the enhanced coupling strength of the system.
-Such increase signifies more frequent and intense scattering processes within the quark-gluon plasma.
+Such increase signifies more frequent and intense scattering processes within the medium.
   2. 2.
@@
 The stronger interactions at higher temperature and with heavier resonances lead to reduced spatial transport.
-Consequently, the meson experiences enhanced confinement and a higher degree of localization within the quark-gluon medium.
+Consequently, the meson experiences enhanced confinement and a higher degree of localization within the medium.
   6. 6.
@@
 The ratios of relaxation times of heavy mesons remain nearly constant over the considered temperature range, indicating a similar temperature response for heavy mesons considered under this study.
-  7. 7.
-    These overall trend point towards the formation of a dense and strongly interacting hadronic medium.
-The enhanced coupling and reduced particle mobility reflects the onset of strong correlations and collective behaviour within the medium.
 ACKNOWLEDGEMENTS